# Normative feeling: socially patterned affective mechanisms

### v3

@@ 1 Introduction
Social norms are an incredibly important aspect of organising a group, as they allow for groups of individuals to coordinate and cooperate with one another. Example norms include, standing a certain distance away in conversation: Too close can be seen as invasive; too far might feel cold or regulating individual greed when managing a shared resource sustainably: too much is seen as selfish but too little is seen as too generous. It has been shown that processes such as punishment and incentives are the key mechanisms underpinning the emergence of norms. They align incentives so that individuals follow norms as to avoid punishment or to gain some reward [5, 36]. However, we know that this isn’t the only reason individuals follow norms, individuals seem to become compelled to follow norms instinctually due to their feelings, even when explicit material punishment is not incurred [16]. For example, psychologist Stanley Milgram instructed his students to ask to take someone’s seat on the subway to examine how people would react to the breaking of a norm. His students attempts where largely unsuccessful, most citing they could not bring themselves to it. Stanley, therefore, decided to try for himself, but when he approached a seated passenger, he found himself stuck. "The words seemed lodged in my trachea and would simply not emerge" , he said. Retreating, he berated himself: ’What kind of craven coward are you?" A few unsuccessful tries later, he managed to blurt out a request. "Taking the man’s seat, I was overwhelmed by the need to behave in a way that would justify my request," he said. "My head sank between my knees, and I could feel my face blanching. I was not role-playing. I actually felt as if I were going to perish." This visceral account of norm breaking suggest that the norms out there in the world and our feelings are interwoven [9]. This shows that concern of how others will perceive us is deeply entrenched in our emotions. The question we want to ask in this article is how did this tight coupling between emotions and norms [26, 35], henceforth social preference, come about evolutionarily?
@@ 1.1 Evolutionary perspective on norms, punishment and mood
Since social norms feature mainly in human societies, many scholars have proposed that these norms arise from cognitive capacities unique to humans. For instance, some theorists argue that social norms are rooted in distinctively human abilities e.g. shared intentionality [42], or a specialized human "norm psychology," [21] shaped by a long history of gene–culture coevolution specific to hominins. Andrighetto et al. argue for a cognitive machinery to detect and reason upon norms that is characterised by a salience mechanism devoted to tracking how much a norm is prominent within a group [5]. More recently, [8] has suggested that social norms may have emerged from cognitive control systems initially selected for the manufacture of complex tools in early hominins [8].
However, these definitions of norms assume that human cognition is essential to norms; meaning we may overlook similar norm-like processes occurring in non-human animals or artificial systems through different cognitive means.
@@
By taking a more inclusive and comparative approach, we may be better able to study the fundamental building blocks of normative cognition. This ultimately, will allow us to better understand their evolutionary emergence, rather than assuming they came about in humans fully formed [13].
Therefore, we see that (minimally) normative processes, such as punishment, occur not only in higher level organisms like humans but in organisms across phyla [36] i.e. simple organisms like insects [46] and bacteria [24]. Bacillus velezensis SQR9 have bacterial free riders who benefit from the collective biofilm but do not contribute to it. They are punished by toxins released by the producers of the biofilm [24], this also occurs in other species through punishment of cheating individuals by cooperator-produced antibiotics [45, 44]. The fact that minimally normative processes such as punishment occur across phyla means:
@@
    We must not only consider normative punishment’s effect on culture but also on the biological fitness of the organisms e.g. for example, free rider phenotypes are selected out genetically but the process is still fundamentally the same as it shapes behaviour, whether through genetics or culture.
@@
    Given normative punishment’ presence across different phyla, it must have emerged very early in evolution or evolved many times.
@@
    This being the case, we must consider normative punishment’ in the evolution of simple and fundamental traits that arose early in evolution e.g. affect (mood, emotions, hormonal processes),
@@
Therefore, using our agent-based model with an evolvable model of mood, we investigate the effect of punishment on the kind of mood mechanism that evolves.
In the process of doing this, we discover a candidate mechanism (mood in the presence of punishment) that results in the emergence of social preferences from individual level selection alone. That is, how a mood mechanism itself can become altruistic via being patterned by normative punishment. Further, we also show how punishment’ allows individuals to align their mood mechanisms in a way so that they can affect each other’s behaviour through affective signals and not physical sanctions, therefore avoiding the costly pitfalls of physical sanctions.
@@ 2.1 Evolution of social preferences
Our attempt to model the normative influence on the evolution/patterning of affect is closely related to several other lines of research. For example with regard to the emergence of social preferences: Frank [16] proposes that in social scenarios such as the Prisoner’s Dilemma some emotions described as "moral sentiments," commit a person to act contrary to their immediate self-interest [16]. For example, the predisposition to feel guilt commits a person to be altruistic, even if cheating were in their material interest. A person with the predisposition to get outraged after having been cheated is committed to punish the cheater, even if it is costly in material terms. Thereby, emotions such as guilt and anger act as "commitment devices" that alter the material incentives. Even so, there should be a material gain from having these emotions, otherwise they would not have evolved in the first place.
Computational studies have investigated how social preferences affect behaviour. Namely by incorporating social preferences by hard coding them into the agent model to see their effect on adaptive group behaviour e.g normative concern motivation [1], equity motivation [12]. Although these studies illuminate how social preferences may affect behaviour in multi-agent settings, they do not explain the emergence of social preferences from purely material concerns i.e. how they come about through individual level selection. Other studies exist that do address the preconditions of internalising social preferences but assume trivial mappings binary between genome and behaviour not based on biological mechanisms [10, 18]. For example, [*]cimpeanu_evolutionary_2025 study the material incentives that may lead to the evolution of guilt but uses a mechanism of guilt that is already functionally pre-specified and linked trivially to a binary genetic character instead of letting the functional relations of the emotion evolve by itself. Another study shows how preferences may be internalised in a psychologically realistic mechanism but this finding relies on high level cognitive capacities found only in humans [6].
In contrast to these other examples, we are interested in how social preference "emerges" and is encoded in our evolvable model of mood out of purely material concern. That is, how individual level selection, in a social setting, patterns a biologically plausible mechanism. In a similar approach to how other research looks at how social dynamics influence the evolution of affect e.g. how predator-pray interactions affect the neuroevolution of affect [19, 23]. However, instead of prey-predator interactions, we want to clarify the effect of normativity, in the form of punishment, on the evolution of mood.
@@ 2.2 Axiomatic Normativity: refining the large space of behavioural regularities
In this section we will explain minimally normative approaches, normative regularities [47] and axiomatic normativity [4]. “Social norms” refer to the rules of a group of people that mark out what is appropriate, allowed, required, or forbidden for various members in different situations. They are typically manifest in common behavioral regularities that are kept in place by social sanctions and social pressure [28].
@@
[47] define “Normative regularities” as “socially maintained patterns of behavioural conformity within a community”. Here "patterns of behavioural conformity" correspond to readily discernible empirical behavioural regularities that support robust predictions and generalisations and whose existence does not depend upon us knowing the underlying mechanisms that produce them. "Social maintenance" here is some form of positive reinforcement of the regularity or negative punishment for breaking away from the regularity. This criterion is needed since not all sources of behavioural conformity are normative. For example, if a group of animals all congregate at the same water source in an otherwise dry landscape, this need not imply that drinking at that water source is a social norm [48]. Therefore, there must be a process of “social maintenance” that maintains the behaviour so it is not purely a consequence of the environment (or as the authors call it environmental scaffolding).
@@
Although this wider minimally cognitive definition is useful in terms of capturing norm like behaviour in non-human animals, it also is very general and does not distinguish between important aspects of normative behaviour. Therefore additional criteria are needed to distinguish between normative and behavioural regularities [4].
@@
There could be social maintenance processes that incentivise individuals in a group to tend to one equilibrium, just as in the case with environmental scaffolding, and others where behaviours tend to many equilibria. For example, in competition, all individuals compete for a resource and are reacting to each other’s behaviour; If an individual exploits a resource more than you, you may have to also exploit it more before they do. One can argue this is social maintenance because individuals are implicitly punishing each other for not being competitive as one another. While competing they and are forced to increase their efforts more and more in order to outcompete others, thus the population will eventually arrive at a maximum possible value of behaviour e.g. peak resource exploitation. Therefore, the outcome would be the same in all populations, only the greediest survive.. This misses out on a key aspect of normativity in our view, which is a degree of arbitrariness [4]. We call processes like competition, social scaffolding i.e. social process which pushes individuals towards the same outcome in every scenario. Another example is prey-predator arms races, where prey will get faster, which in turn forces the predator to be faster, but will ultimately tend to one equilibrium - the fastest possible speed.
@@
Therefore, we introduce a multi-equilibria criterion which can distinguish between those two types of normative regularities. In general, we suggest the axiomatic norm definition, an approach where norms are defined by a set of criteria which is as follows:
@@
This norm definition provides a lens to distinguish between different social processes i.e. processes that result in one behavioural equilibrium and processes where many equilibria occur, acquiring a hallmark of culture and normative behaviour: arbitrariness in behaviour between groups. This allows us to talk about the wide space of behavioural regularities in a more granular fashion.
The definition also allows for the study of punishment/incentivisation on behaviour whether its genetically or culturally encoded. This allows us to also examine the effect of punishment on genetic evolution as well as on cultural evolution. This is an important aspect of the definition as we believe how normative processes shape biological behaviour is understudied.
@@
2.3 Normative punishment and it’s impact on the evolution of mood
@@
There could be several candidate mechanisms that maintain multiple equilibria normative regularities, but in this paper we will investigate normative punishment. We define punishment as an action an agent takes that negatively affects another agent and is conditionally triggered by that same agent violating an explicit criterion e.g. taking too much resource. This negatively effects the agent’s evolutionary success and therefore the propagation for their genetic or cultural strategy based on this criterion from the punishing agent. For example,the negative effect could be decreasing their chances of survival or ability to reproduce (so they do not propagate genetically), or their ability to spread their cultural profile such as diminishing their status so they are not copied by others (thus limiting the spread of their behaviour culturally).
Using the above definition of punishment, we can see that punishment does occur not only in higher level organisms like humans, but across phyla [36], even in simple organisms like insects [46] and bacteria [24]. For example, e.g bacteria reduce the evolutionary success of free riders by releasing antibiotics, therefore killing and impeding free riders from passing on their behavioural phenotype to the next generation. We can now assert punishment, according to our definition, is present in many species and is ergo, widely conserved in evolution (or occurred many time). Therefore, its influence must be considered in an evolutionary account of mood.
In our experiments, we will verify if punishment (under our definition) results in multiple behavioural equilibria, therefore making it qualitatively different to other social processes e.g. competition, and in our view, truly normative . We will then compare how mood evolves and emerges under punishment i.e. explicit regulation of anothers’ fitness and other non multi-equilibria social process’, i.e. competition - implicit regulation of anothers’ fitness and compare the effect of these processes on the evolution of the mood mechanisms.
@@ 3.1 Model overview
We use an evolutionary multi-agent simulation with agents who have evolvable mood mechanisms that modulate their behaviour. Agents (N=100) take turns to consume a dynamically replenishing resource (eat step) and punish (sanction step). This is because this is a space-less simulation with no topology, so agents punish a random subset of the agents by subtracting from their energy. Energy level will determine procreation, and thus determine what agent parameters become more prevalent in the society. After initialisation, the simulation proceeds in rounds. Each round has a different, randomised order of all agents, and each of the following steps is performed in the following order for 2,000 time steps. Each condition is repeated 1000 times and metrics averaged over runs. And, at the start of each simulation, all agent values including internal values and weights of the mood mechanism are randomised.
@@ 3.5 Death and Reproduction
During the reproduction step, all agents with energy below 0 are removed. Then any agent with an energy level larger than 10 gets to reproduce. Reproduction means that we generate a copy of the agent with the same traits, mutated with a $0.1$ chance (this is a global parameter that applies to all traits being mutated). Mutations are adding Gaussian noise with 0 mean and $1$ variance to each of the agents parameters. The energy of the child and parent are both set at half the parent’s prior energy level. Further, to ensure evolution does not stop when populations are stable, we assume at each simulation step that each agent has a 1/100 chance of dying. Therefore, this means that each agent has on average a 100 time-step life span instead of being immortal, allowing for other agents to be born, and therefore allowing variation through mutation.

### v4

@@ 1 Introduction
Social norms are an incredibly important aspect of organising a group, as they allow for groups of individuals to coordinate and cooperate with one another. Example norms include, standing a certain distance away in conversation: Too close can be seen as invasive; too far might feel cold or regulating individual greed when managing a shared resource sustainably: too much is seen as selfish but too little is seen as too generous. It has been shown that processes such as punishment and incentives are the key mechanisms underpinning the emergence of norms. They align incentives so that individuals follow norms as to avoid punishment or to gain some reward [5, 36]. However, we know that this isn’t the only reason individuals follow norms; individuals seem to become compelled to follow norms instinctually due to their feelings, even when explicit material punishment is not incurred [16]. For example, psychologist Stanley Milgram instructed his students to ask to take someone’s seat on the subway to examine how people would react to the breaking of a norm. His students’ attempts were largely unsuccessful, most citing they could not bring themselves to it. Stanley, therefore, decided to try for himself, but when he approached a seated passenger, he found himself stuck. "The words seemed lodged in my trachea and would simply not emerge", he said. Retreating, he berated himself: ’What kind of craven coward are you?" A few unsuccessful tries later, he managed to blurt out a request. "Taking the man’s seat, I was overwhelmed by the need to behave in a way that would justify my request," he said. "My head sank between my knees, and I could feel my face blanching. I was not role-playing. I actually felt as if I were going to perish." This visceral account of norm breaking suggest that the norms out there in the world and our feelings are interwoven [9]. This shows that the concern of how others will perceive us is deeply entrenched in our emotions. The question we want to ask in this article is how this tight coupling between emotions and norms [26, 35], henceforth social preference, come about evolutionarily?
@@ 1.1 Evolutionary perspective on norms, punishment and mood
Since social norms feature mainly in human societies, many scholars have proposed that these norms arise from cognitive capacities unique to humans. For instance, some theorists argue that social norms are rooted in distinctively human abilities e.g. shared intentionality [42], or a specialised human "norm psychology," [21] shaped by a long history of gene–culture coevolution specific to hominins. Andrighetto et al. argue for a cognitive machinery to detect and reason upon norms that is characterised by a salience mechanism devoted to tracking how much a norm is prominent within a group [5]. More recently, [8] has suggested that social norms may have emerged from cognitive control systems initially selected for the manufacture of complex tools in early hominins [8].
However, these definitions of norms assume that human cognition is essential to norms, meaning we may overlook similar norm-like processes occurring in non-human animals or artificial systems through different cognitive means.
@@
By taking a more inclusive and comparative approach, we may be better able to study the fundamental building blocks of normative cognition. This ultimately will allow us to better understand their evolutionary emergence, rather than assuming they came about in humans fully formed [13].
Therefore, we see that (minimally) normative processes, such as punishment, occur not only in higher-level organisms like humans but in organisms across phyla [36], i.e. simple organisms like insects [46] and bacteria [24]. Bacillus velezensis SQR9 have bacterial free riders who benefit from the collective biofilm but do not contribute to it. They are punished by toxins released by the producers of the biofilm [24]. This also occurs in other species through punishment of cheating individuals by cooperator-produced antibiotics [45, 44]. The fact that minimally normative processes such as punishment occur across phyla means:
@@
    We must not only consider normative punishment’s effect on culture but also on the biological fitness of the organisms, e.g. for example, free rider phenotypes are selected out genetically, but the process is still fundamentally the same as it shapes behaviour, whether through genetics or culture.
@@
    Given normative punishment’s presence across different phyla, it must have emerged very early in evolution or evolved many times.
@@
    This being the case, we must consider normative punishment in the evolution of simple and fundamental traits that arose early in evolution, e.g. affect (mood, emotions, hormonal processes),
@@
Therefore, using our agent-based model with an evolvable mood mechanism, we investigate the effect of punishment on the kind of mood mechanism that evolves.
In the process of doing this, we discover a candidate mechanism (mood in the presence of punishment) that results in the emergence of social preferences from individual-level selection alone. That is, how a mood mechanism itself can become altruistic via being patterned by normative punishment. Further, we also show how punishment allows individuals to align their mood mechanisms in a way so that they can affect each other’s behaviour through affective signals and not physical sanctions, therefore avoiding the costly pitfalls of physical sanctions.
@@ 2.1 Evolution of social preferences
Our attempt to model the normative influence on the evolution/patterning of affect is closely related to several other lines of research. For example, with regard to the emergence of social preferences, Frank [16] proposes that in social scenarios such as the Prisoner’s Dilemma, some emotions described as "moral sentiments" commit a person to act contrary to their immediate self-interest [16]. For example, the predisposition to feel guilt commits a person to be altruistic, even if cheating were in their material interest. A person with the predisposition to get outraged after having been cheated is committed to punish the cheater, even if it is costly in material terms. Thereby, emotions such as guilt and anger act as "commitment devices" that alter the material incentives. Even so, there should be a material gain from having these emotions, otherwise they would not have evolved in the first place.
Computational studies have investigated how social preferences affect behaviour. Namely, by incorporating social preferences by hard-coding them into the agent model to see their effect on adaptive group behaviour e.g normative concern motivation [1], equity motivation [12]. Although these studies illuminate how social preferences may affect behaviour in multi-agent settings, they do not explain the emergence of social preferences from purely material concerns, i.e. how they come about through individual-level selection. Other studies exist that do address the preconditions of internalising social preferences but assume trivial mappings binary between genome and behaviour, not based on biological mechanisms [10, 18]. For example, [*]cimpeanu_evolutionary_2025 study the material incentives that may lead to the evolution of guilt but uses a mechanism of guilt that is already functionally pre-specified and linked trivially to a binary genetic character instead of letting the functional relations of the emotion evolve by itself. Another study shows how preferences may be internalised in a psychologically realistic mechanism, but this finding relies on high-level cognitive capacities found only in humans [6].
In contrast to these other examples, we are interested in how social preference "emerges" and is encoded in our evolvable model of mood out of purely material concern. That is, how individual-level selection, in a social setting, patterns a biologically plausible mechanism. In a similar approach to how other research looks at how social dynamics influence the evolution of affect e.g. how predator-prey interactions affect the neuroevolution of affect [19, 23]. However, instead of prey-predator interactions, we want to clarify the effect of normativity, in the form of punishment, on the evolution of mood.
@@ 2.2 Axiomatic Normativity: refining the large space of behavioural regularities
In this section, we will explain minimally normative approaches, normative regularities [47] and axiomatic normativity [4]. “Social norms” refer to the rules of a group of people that mark out what is appropriate, allowed, required, or forbidden for various members in different situations. They are typically manifest in common behavioural regularities that are kept in place by social sanctions and social pressure [28].
@@
[47] define “Normative regularities” as “socially maintained patterns of behavioural conformity within a community”. Here, "patterns of behavioural conformity" correspond to readily discernible empirical behavioural regularities that support robust predictions and generalisations and whose existence does not depend upon us knowing the underlying mechanisms that produce them. "Social maintenance" here is some form of positive reinforcement of the regularity or negative punishment for breaking away from the regularity. This criterion is needed since not all sources of behavioural conformity are normative. For example, if a group of animals all congregate at the same water source in an otherwise dry landscape, this need not imply that drinking at that water source is a social norm [48]. Therefore, there must be a process of “social maintenance” that maintains the behaviour so it is not purely a consequence of the environment (or as the authors call it environmental scaffolding).
@@
Although this wider minimally cognitive definition is useful in terms of capturing norm-like behaviour in non-human animals, it is also very general and does not distinguish between important aspects of normative behaviour. Therefore, additional criteria are needed to distinguish between normative and behavioural regularities [4].
@@
There could be social maintenance processes that incentivise individuals in a group to tend to one equilibrium, just as in the case with environmental scaffolding, and others where behaviours tend to many equilibria. For example, in competition, all individuals compete for a resource and are reacting to each other’s behaviour; If an individual exploits a resource more than you, you may have to also exploit it more before they do. One can argue that this is social maintenance because individuals are implicitly punishing each other for not being as competitive as one another. While competing, they are forced to increase their efforts more and more in order to outcompete others, thus the population will eventually arrive at a maximum possible value of behaviour, e.g. peak resource exploitation. Therefore, the outcome would be the same in all populations; only the greediest survive. This misses out on a key aspect of normativity in our view, which is a degree of arbitrariness [4]. We call processes like competition, social scaffolding, i.e. social process which pushes individuals towards the same outcome in every scenario. Another example is prey-predator arms races, where prey will get faster, which in turn forces the predator to be faster, but will ultimately tend to one equilibrium - the fastest possible speed.
@@
Therefore, we introduce a multi-equilibria criterion which can distinguish between those two types of normative regularities. In general, we suggest the axiomatic norm definition, an approach where norms are defined by a set of criteria, which is as follows:
@@
This norm definition provides a lens to distinguish between different social processes, i.e. processes that result in one behavioural equilibrium and processes where many equilibria occur, acquiring a hallmark of culture and normative behaviour: arbitrariness in behaviour between groups. This allows us to talk about the wide space of behavioural regularities in a more granular fashion.
The definition also allows for the study of punishment/incentivisation on behaviour, whether it is genetically or culturally encoded. This allows us to also examine the effect of punishment on genetic evolution as well as on cultural evolution. This is an important aspect of the definition, as we believe how normative processes shape biological behaviour is understudied.
@@
2.3 Normative punishment and its impact on the evolution of mood
@@
There could be several candidate mechanisms that maintain multiple equilibria normative regularities, but in this paper, we will investigate normative punishment. We define punishment as an action an agent takes that negatively affects another agent and is conditionally triggered by that same agent violating an explicit criterion, e.g. taking too much resource. This negatively affects the agent’s evolutionary success and, therefore, the propagation of their genetic or cultural strategy based on this criterion from the punishing agent. For example, the negative effect could be decreasing their chances of survival or ability to reproduce (so they do not propagate genetically), or their ability to spread their cultural profile, such as diminishing their status so they are not copied by others (thus limiting the spread of their behaviour culturally).
Using the above definition of punishment, we can see that punishment does occur not only in higher-level organisms like humans, but across phyla [36], even in simple organisms like insects [46] and bacteria [24]. For example, e.g bacteria reduce the evolutionary success of free riders by releasing antibiotics, therefore killing and impeding free riders from passing on their behavioural phenotype to the next generation. We can now assert that punishment, according to our definition, is present in many species and is ergo, widely conserved in evolution (or occurred many times). Therefore, its influence must be considered in an evolutionary account of mood.
In our experiments, we will verify if punishment (under our definition) results in multiple behavioural equilibria, therefore making it qualitatively different to other social processes, e.g. competition, and in our view, truly normative. We will then compare how mood evolves and emerges under punishment, i.e. explicit regulation of anothers’ fitness and other non-multi-equilibria social processes, i.e. competition - implicit regulation of anothers’ fitness and compare the effect of these processes on the evolution of the mood mechanisms.
@@ 3.1 Model overview
We use an evolutionary multi-agent simulation with agents who have evolvable mood mechanisms that modulate their behaviour. Agents (N=100) take turns to consume a dynamically replenishing resource (eat step) and punish (sanction step). This is because this is a spaceless simulation with no topology, so agents punish a random subset of the agents by subtracting from their energy. Energy level will determine procreation, and thus determine which agent parameters become more prevalent in the population. After initialisation, the simulation proceeds in rounds. Each round has a different, randomised order of all agents, and each of the following steps is performed in the following order for 2,000 time steps. Each condition is repeated 1000 times, and metrics are averaged over runs. And, at the start of each simulation, all agent values, including internal values and weights of the mood mechanism, are randomised.
@@ 3.5 Death and Reproduction
During the reproduction step, all agents with energy below 0 are removed. Then any agent with an energy level larger than 10 gets to reproduce. Reproduction means that we generate a copy of the agent with the same traits, mutated with a $0.1$ chance (this is a global parameter that applies to all traits being mutated). Mutations are adding Gaussian noise with 0 mean and $1$ variance to each of the agents’ parameters. The energy of the child and parent are both set at half the parent’s prior energy level. Further, to ensure evolution does not stop when populations are stable, we assume at each simulation step that each agent has a 1/100 chance of dying. Therefore, this means that each agent has, on average, a 100-time-step life span instead of being immortal, allowing for other agents to be born, and therefore allowing variation through mutation.